OpenStudy (anonymous):

Norma and David crawled to the barn and then hopped back to the house. They crawled at 300 centimeters per minute and hopped 400 centimeters per minute. If the round trip took 7 minutes, how long did they crawl? How far was it to the barn?

OpenStudy (anonymous):

300t = 400 ( 7 -t) 7t = 28 ->t = 28/7 = 4 min The barn is 300 * 4 = 1200 cm = 12m away

OpenStudy (anonymous):

But if the round trip was a total of 7 minutes, how was the barn 12m away?

OpenStudy (anonymous):

Did you see 1200 cm = 12m?

OpenStudy (anonymous):

Oh! Sorry.